OpenStudy (anonymous):

If a and b are positive numbers, find the maximum value of f(x)=xa(1−x)b, 0≤x≤1 Your answer may depend on a and b.

OpenStudy (amoodarya):

f(x)=ab x(1-x) f'=ab (1-2x)=0 then x=0.5 f(0)=0 f(1)=0 f(0.5)=ab (0.25) so if ab>0 then max is f(.5)=0.25 ab if ab <0 max f(x)=0

OpenStudy (amoodarya):